OpenStudy (ashley1nonly):

Carbon-14 has a half-life of 5720 years and this is a first order reaction. If a piece of wood has converted 25% of the carbon-14, then how old is it? A) 11440 years B) 2375 years C) 4750 years D) 4290 years E) 1430 years Answer: B ) For a particular first-order reaction, it takes 24 minutes for the concentration of the reactant to decrease to 25% of its initial value. What is the value for rate constant (in s-1) for the reaction? A) 2.0 × 10-4 s-1 B) 9.6 × 10-4 s-1 C) 1.2 × 10-2 s-1 D) 5.8 × 10-2 s-1 Answer: B Diff: 3 Page Ref: 13.4

OpenStudy (ashley1nonly):

ok i know how to solve the problem but i am having problems with knowning when do i subtract from 100

OpenStudy (anonymous):

what was the equation again t1/2=log2/k ?

OpenStudy (ashley1nonly):

the first one i would do 100-25 the second i would just do ln[25]=-kt+ln[100]

OpenStudy (ashley1nonly):

t1/2= .693/k

OpenStudy (anonymous):

kk thx let me look at this

OpenStudy (ashley1nonly):

first one ln[75]=-kt+ln[100]

OpenStudy (anonymous):

would you divide 25/100 for the 1st one?

OpenStudy (ashley1nonly):

the answer wouldnt be correct you have to do 100-25=75 which confuses me, when do you subtract from 100 and when wouldnt you

OpenStudy (anonymous):

I believe since they gave you half life you can subtract from 100, and the 2nd question you do not have half life given

OpenStudy (anonymous):

I forgot why, but if you have half life given you can subtract 100

OpenStudy (ashley1nonly):

ok i have another question

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

I got it since carbon 12natural is about 99%, and we are using carbon 14, we subtract 100-25

OpenStudy (ashley1nonly):

???

OpenStudy (ashley1nonly):

oh i remeber when i solved for that in 1045 ok.

OpenStudy (anonymous):

kk

OpenStudy (ashley1nonly):

so when i have a half life is when i will subtract from 100 basically

OpenStudy (anonymous):

yes when its says carbon 14 then subtract 100

OpenStudy (ashley1nonly):

ok i have two more questions and i just posted one of them

OpenStudy (anonymous):

kk